# O. Ragniscoa, F.Zullob

### v1

@@ 5.3.1 The numerics of the continuous case
In order to numerically integrate the equation (5.34), we set $\alpha=1$ and $C=1$. As for the discrete case, we plot three orbits. The orbits correspond to the following initial conditions: $(Q(0),P(0))=(2,0)$, $(Q(0),P(0))=(5,0)$ and (Q(0),P(0))=(7,5)(Q(0),P(0))=(-7,5). The corresponding values of the energies are $6.61584$, $77.80793$ and $456.00834$. The plots on the right in Figure (3), in blue, are the plots of the analytical curves corresponding to the level curves defined by the Hamiltonian in these three cases and coincide with the orbits. The curve in red is the curve defined by $\dot{P}=0$: the equilibrium position is where this curve intersect the $Q$ axis.
Figure 3: Plot of three orbits of the system (5.34): on the left there is the numeric, on the right the plot of the corresponding Hamiltonian level curves. In red the curve defined by $\dot{P}=0$.
@@
As specified in Remark (4.7), if the functions $\eta(x)$ and $\kappa(x)$ are in $L^{1}$, the equations of motion (4.58) still give a finite result. Let us take the following functions
$\eta(x)=\frac{2\sigma}{\pi(1+x^{2})},\quad\sigma^{2}=1,\quad\kappa(x)=\frac{2}{\pi}\frac{C\text{sgn}(x)}{(1+x^{2})}.$ (5.37)
In order to deal with the equation for $\dot{P}$ in (4.58), we use the following result:
$\int_{0}^{\infty}\frac{1}{1+x^{2}}\exp(\frac{y}{1+x^{2}})dx=\frac{\pi}{2}\exp(y/2)I_{0}(y/2),$ (5.38)
@@
$\frac{1}{\pi}\int_{0}^{\infty}\exp(z\cos(\xi))d\xi=I_{0}(z).$ (5.39)
Indeed, by the change of variables $x=\tan(\xi/2)$ (5.38) becomes
$\int_{0}^{\infty}\frac{1}{1+x^{2}}\exp(\frac{y}{1+x^{2}})dx=\frac{1}{2}\int_{0}^{\pi}\exp\left(\frac{y}{2}(1+\cos(\xi))\right)d\xi$ (5.40)
and formula (5.38) follows. From the definitions (5.37) and fromula (5.38) we get for the equations of motion (4.58)
$\dot{Q}=P,\quad\dot{P}=1-\sigma\left(e^{\frac{\sigma Q+C}{\pi}}I_{0}\left(\frac{\sigma Q+C}{\pi}\right)+e^{\frac{\sigma Q-C}{\pi}}I_{0}\left(\frac{\sigma Q-C}{\pi}\right)\right)$ (5.41)
@@ 6 Concluding remarks
In this paper we proposed an extension of the integrable Volterra model to the case of infinitely many species, either countable or uncountable. The corresponding models have been shown to possess the property of being maximally superintegrable. In our opinion, though the generalizations proposed are non trivial, they do not lead to an enrichment of the dynamics of the system. Indeed, within the class of sequences or continuous functions we have considered, in the infinite limit, both discrete and continuous, the behaviour of the system looks even simpler than the one characterizing the case of a finite number of species. We may assert that the maximal superintegrability, already discovered in [17], gets in a sense strengthened by the infinite limit, inasmuch as there exists a general formula showing that, in canonical coordinates, the Hamiltonian takes, in way unexpectedly, the classical one body feature ${\mathcal{H}}=\frac{P^{2}}{2}+V(Q)$. Moreover, the potential $V(Q)$ supports just one equilibrium position in the general case, around which the motion foliates in a family of closed orbits and represents an infinite potential well. In a given class of special cases one can have open orbits. These results are strongly related to the family of sequences for the growth coefficients $\epsilon_{n}$ or the family of growth function $\epsilon(x)$ chosen. In principle, a different choice of these coefficients or function can destroy the simple Hamiltonian structure described in this work. The instances here given are, in our opinion, significant for the following reasons: they give, in the infinite limit, both in the discrete and in the continuous case, a finite and specific Hamiltonian model. Also, the uncountable case can be seen as the natural limit of the denumerable one. Further, the sequences $\eta_{n}$ and $\kappa_{n}$, or the functions $\eta(x)$ and $\kappa(x)$ for the continuous case, are essentially arbitrary (apart from the fact of belonging to a suitable summable or integrable space) and the properties of the corresponding Hamiltonian models are universal, in the sense that, fixed the sequences $\eta_{n}$ and $\kappa_{n}$ or the functions $\eta(x)$ and $\kappa(x)$, the behavior of the orbits are fixed (or periodic or open for specific cases). Finally, the connection between the growth coefficients and the automatic sequences, like the Rudin-Shapiro sequence, briefly mentioned in the text, is somehow unexpected and surely deserves much attention in future works. As for the unanswered questions: we have not been able to find an explicit formula for the period of the orbits, apart the linearized case. For a finite number of species we have shown [18] that, in suitable cases, the period of the linearized system is a very good approximation of the period of the orbits of the non linearized model. We expect that, if the nonlinearity is not too strong, the observation remains true also in our examples, but we didn’t give any quantitative statement about this. Finally, it would be interesting to expand the class of sequences or functions for $\epsilon_{n}$ and $\epsilon(x)$ here given to other families: as said, the Hamiltonian structure that we obtained may be destroyed and richer models, for example with a greater number of equilibrium positions and different type of behaviors for the orbits, could be found.

### v2

@@ 5.3.1 The numerics of the continuous case
In order to numerically integrate the equation (161), we set $\alpha=1$ and $C=1$. As for the discrete case, we plot three orbits. The orbits correspond to the following initial conditions: $(Q(0),P(0))=(2,0)$, $(Q(0),P(0))=(5,0)$ and (Q(0),P(0))=(7,5)(Q(0),P(0))=(-7,5). The corresponding values of the energies are $6.61584$, $77.80793$ and $456.00834$. The plots on the right in Figure (3), in blue, are the plots of the analytical curves corresponding to the level curves defined by the Hamiltonian in these three cases and coincide with the orbits. The curve in red is the curve defined by $\dot{P}=0$: the equilibrium position is where this curve intersect the $Q$ axis.
Figure 3: Plot of three orbits of the system (161): on the left there is the numeric, on the right the plot of the corresponding Hamiltonian level curves. In red the nullclines defined by $\dot{P}=0$ and $\dot{Q}=0$.
@@
As specified in Remark (4.7), if the functions $\eta(x)$ and $\kappa(x)$ are in $L^{1}$, the equations of motion (124) still give a finite result. Let us take the following functions
$\eta(x)=\frac{2\sigma}{\pi(1+x^{2})},\quad\sigma^{2}=1,\quad\kappa(x)=\frac{2}{\pi}\frac{C\text{sgn}(x)}{(1+x^{2})}.$ (164)
In order to deal with the equation for $\dot{P}$ in (124), we use the following result:
$\int_{0}^{\infty}\frac{1}{1+x^{2}}\exp(\frac{y}{1+x^{2}})dx=\frac{\pi}{2}\exp(y/2)I_{0}(y/2),$ (165)
@@
$\frac{1}{\pi}\int_{0}^{\infty}\exp(z\cos(\xi))d\xi=I_{0}(z).$ (166)
Indeed, by the change of variables $x=\tan(\xi/2)$ (165) becomes
$\int_{0}^{\infty}\frac{1}{1+x^{2}}\exp(\frac{y}{1+x^{2}})dx=\frac{1}{2}\int_{0}^{\pi}\exp\left(\frac{y}{2}(1+\cos(\xi))\right)d\xi$ (167)
and formula (165) follows. From the definitions (164) and formula (165) we get for the equations of motion (124)
$\dot{Q}=P,\quad\dot{P}=1-\sigma\left(e^{\frac{\sigma Q+C}{\pi}}I_{0}\left(\frac{\sigma Q+C}{\pi}\right)+e^{\frac{\sigma Q-C}{\pi}}I_{0}\left(\frac{\sigma Q-C}{\pi}\right)\right)$ (168)
@@ 6 Concluding remarks
In this paper we proposed an extension of the integrable Volterra model to the case of infinitely many species, either countable or uncountable. The corresponding models have been shown to possess the property of being maximally superintegrable since reduce to a system with just one degree of freedom. This property is independent on the choice of the parameters, but is essentially related to the fact that, in the integrable case, the interaction matrix $A$ has always rank 2. The one dimensional equations of motion are (16)-(17) for the discrete case and (71)-(72) for the continuous case. Clearly, it is necessary that the sequences $\epsilon_{n}$, $\eta_{n}$ and $\kappa_{n}$ for the discrete case or the functions $\epsilon(x)$, $\eta(x)$ and $\kappa(x)$ for the continuous case are such that the corresponding equations converge.
In our opinion, though the generalizations proposed are non trivial, they do not lead to an enrichment of the dynamics of the system. Indeed, within the class of sequences or continuous functions we have considered, in the infinite limit, both discrete and continuous, the behaviour of the system looks even simpler than the one characterizing the case of a finite number of species. We may assert that the maximal superintegrability, already discovered in [16], gets in a sense strengthened by the infinite limit, inasmuch as there exists a general formula showing that, in canonical coordinates, the Hamiltonian takes, in a way unexpectedly, the classical one body feature ${\mathcal{H}}=\frac{P^{2}}{2}+V(Q)$. For the discrete case this is true if the sequence $\epsilon_{n}$ is chosen like in (32) or (50), i.e. is a set of +1+1 and 1-1, normalized in $\ell^{2}$, distributed evenly or oddly on the real line, whereas for the continuous case we take a normalized step-like piecewise constant function, evenly or oddly distributed on the real line. In order to get the convergence, the sequences $\eta_{n}$ and $\kappa_{n}$ can belong to the set of rapidly decreasing sequences $S^{\infty}$ or to the larger space $\ell^{1}$, and the corresponding equivalent assumptions must be made about the functions $\eta(x)$ and $\kappa(x)$ in the continuous case.
The explicit form the potential $V(Q)$ takes depends on the particular choices of the sequences $\eta_{n}$, $\kappa_{n}$ (or $\eta(x)$ and $\kappa(x)$) for the continuous case. We give some examples. Also, we have shown that with our particular choices of the growth rates $\epsilon_{n}$ ($\epsilon(x)$), the Hamiltonian system ${\mathcal{H}}=\frac{P^{2}}{2}+V(Q)$ possesses bounded periodic orbits if the sequence $\eta_{n}$ has not a fixed sign or if $\eta_{n}\geq 0$ for all $n$, whereas possesses open orbits if $\eta_{n}\leq 0$ for all $n$. Examples of all these cases have been given. The same statements are true in the continuous case by requiring respectively that the function $\eta(x)$ has not a fixed sign, is non-positive or is non-negative.
We have also shown that taking an even or odd Kernel (or growth rates) is not limitative since it is possible to take a combination of the even and odd Hamiltonian flow and get a mixed Hamiltonian. The properties of the corresponding orbits are also been given.
The above results are strongly related to the family of sequences for the growth coefficients $\epsilon_{n}$ or the family of growth function $\epsilon(x)$ chosen. In principle, a different choice of these coefficients or function can destroy the simple Hamiltonian structure described in this work. The instances here given are, in our opinion, significant for the following reasons: they give, in the infinite limit, both in the discrete and in the continuous case, a finite and specific Hamiltonian model. Also, the uncountable case can be seen as the natural limit of the denumerable one. Further, the sequences $\eta_{n}$ and $\kappa_{n}$, or the functions $\eta(x)$ and $\kappa(x)$ for the continuous case, are essentially arbitrary (apart from the fact of belonging to a suitable summable or integrable space) and the properties of the corresponding Hamiltonian models are universal, in the sense that, fixed the sequences $\eta_{n}$ and $\kappa_{n}$ or the functions $\eta(x)$ and $\kappa(x)$, the behaviour of the orbits are fixed (either periodic or open for specific cases).
Finally, the connection between the growth coefficients and the automatic sequences, like the Rudin-Shapiro sequence, briefly mentioned in the text, is somehow unexpected and surely deserves much attention in future works.
As for the unanswered questions: we have not been able to find an explicit formula for the period of the orbits, apart for the linearized case. For a finite number of species we have shown [17] that, in suitable cases, the period of the linearized system is a very good approximation of the period of the orbits of the non linearized model. We expect that, if the nonlinearity is not too strong, the observation remains true also in our examples, but we didn’t give any quantitative statement about this. Finally, it would be interesting to expand the class of sequences or functions for $\epsilon_{n}$ and $\epsilon(x)$ here given to other families: as said, the Hamiltonian structure that we obtained may be destroyed and richer models, for example with a greater number of equilibrium positions and different type of behaviours for the orbits, could be found.